# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ V Privacy analysis
$\textstyle\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{p})\tilde{z}_{i}^{t}-\nabla l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{\prime p})\tilde{z}_{i}^{\prime t}\|_{1}$ (10)
$\textstyle\leq\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{p})\|_{2}\|\tilde{z}_{i}^{t}-\tilde{z}_{i}^{\prime t}\|_{1}$
$\textstyle\quad+\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\sum_{p\neq k}^{t}\|\nabla l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{p})-\nabla l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{p})\|_{2}\|\tilde{z}_{i}^{\prime t}\|_{2}$
$\textstyle\quad+\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\|\nabla l(R_{i}\boldsymbol{x}_{i}^{t};\xi_{i}^{k})-\nabla l(R_{i}\boldsymbol{x}_{i}^{\prime t};\xi_{i}^{\prime k})\|_{2}\|\tilde{z}_{i}^{\prime t}\|_{2}$
$\textstyle\leq\frac{\sqrt{n_{i}}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}}(\Delta_{i,z}^{t+1}+\Delta_{i,z}^{t})\!+\!\frac{\sqrt{n_{i}}\bar{L}_{l}d_{z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\Delta_{i,x}^{t}\!+\!\frac{2\sqrt{n_{i}}L_{l}d_{z}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)},$
where we have used the fact that the convergence of Algorithm 1 ensures $\max_{i\in[m],t>0}\{\|z_{i}^{t}\|_{2}\}\leq d_{z}$ for some $d_{z}>0$.
Substituting (9) and (10) into (8), we arrive at
$\textstyle\Delta_{i,x}^{t+1}\leq\left(1-\bar{w}+\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{t}+\frac{\sqrt{n_{i}}\bar{L}_{l}d_{z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\right)\Delta_{i,x}^{t}$
@@
$\textstyle\quad+\frac{2\sqrt{n_{i}}L_{h}\lambda_{x}^{t}}{t+1}+\frac{2\sqrt{n_{i}}d_{z}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)}.$ (11)
@@
$\|z_{i}^{t+1}-z_{i}^{\prime t+1}\|_{1}\leq(1-\bar{w})\|z_{i}^{t}-z_{i}^{\prime t}\|_{1}$
$\textstyle+\frac{\sqrt{r}\lambda_{z}^{t}}{t+1}\sum_{p\neq k}^{t}\|\nabla_{y}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})\!-\!\nabla_{y}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{p})\|_{2}$
$\textstyle+\frac{\sqrt{r}\lambda_{z}^{t}}{t+1}\|\nabla_{y}h(R_{i}\boldsymbol{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{k})\!-\!\nabla_{y}h(R_{i}\boldsymbol{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime k})\|_{2}.$
@@
$\Delta_{i,z}^{t+1}\textstyle\leq(1-\bar{w})\Delta_{i,z}^{t}+\sqrt{r}\bar{L}_{h}\lambda_{z}^{t}\Delta_{i,x}^{t}$ (12)
@@
$\textstyle\leq\left(1\!-\!\bar{w}\!+\!\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{t}+\frac{\sqrt{n_{i}}\bar{L}_{l}d_{z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\!+\!\sqrt{r}(L_{l}\lambda_{y}^{t}\!+\!\bar{L}_{h}\lambda_{z}^{t})\right)\Delta_{i,x}^{t}$
@@
$\textstyle+\frac{2d_{l}\lambda_{y}^{t}}{t+1}+\frac{2\sqrt{n_{i}}L_{h}\lambda_{x}^{t}}{t+1}+\frac{2\sqrt{n_{i}}d_{z}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)}+\frac{2\sqrt{r}L_{h}\lambda_{z}^{t}}{t+1}.$ (13)
Since the decaying rates of stepsizes $\lambda_{x}^{t}$, $\lambda_{y}^{t}$, and $\lambda_{z}^{t}$ satisfy $01\!>\!v_{x}\!>\!v_{z}\!>\!v_{y}\!>\!$ 0, we can choose proper initial stepsizes such that the following inequality always holds:
$\textstyle\Delta_{i,x}^{t+1}+\Delta_{i,y}^{t+1}+\Delta_{i,z}^{t+1}$ $\textstyle\leq\left(1-\frac{\bar{w}}{2}\right)\left(\Delta_{i,x}^{t}+\Delta_{i,y}^{t}+\Delta_{i,z}^{t}\right)$
$\textstyle\quad+\frac{C_{0}}{(t+1)^{\min\{1+v_{x}-v_{z},1+v_{y}\}}},$ (14)
where $C_{0}\!=\!2(d_{l}\lambda_{y}^{0}+\sqrt{n_{i}}\lambda_{x}^{0}(L_{h}+d_{z}L_{l}(\lambda_{z}^{0})^{-1})+\sqrt{r}L_{h}\lambda_{z}^{0})$.
By applying Lemma 11 in [47] to (14) and using the relation $\Delta_{i,x}^{0}\!=\!\Delta_{i,y}^{0}\!=\!\Delta_{i,z}^{0}\!=\!0$, we have
$\Delta_{i,x}^{t}+\Delta_{i,y}^{t}+\Delta_{i,z}^{t}\leq\frac{C_{1}}{(t+1)^{\min\{1+v_{x}-v_{z},1+v_{y}\}}},$ (15)
with $C_{1}\!=\!\frac{4C_{0}}{\bar{w}}(\frac{4\min\{1+v_{x}-v_{z},1+v_{y}\}}{e\ln(\frac{4}{4-\bar{w}})})^{\min\{1+v_{x}-v_{z},1+v_{y}\}}$.
Substituting $\Delta_{i,x}^{t}\leq C_{1}(t+1)^{-\min\{1+v_{x}-v_{z},1+v_{y}\}}$ into (7) and using Lemma 11 in [47], we obtain
$\vspace{-0.2em}\textstyle\Delta_{i,y}^{t}\leq\frac{C_{y}}{(t+1)^{1+v_{y}}}~{}\text{with}~{}C_{y}=\frac{2C_{2}}{\bar{w}}(\frac{4(1+v_{y})}{e\ln(\frac{2}{2-\bar{w}})})^{1+v_{y}},\vspace{-0.2em}$ (16)
where $C_{2}=(C_{1}L_{l}\sqrt{r}+2d_{l})\lambda_{y}^{0}$ and $C_{1}$ is given in (15).
Substituting $\Delta_{i,z}^{t}\!\leq\!C_{1}(t+1)^{-\min\{1+v_{x}-v_{z},1+v_{y}\}}$ and (16) into (11), we have
$\vspace{-0.2em}\textstyle\Delta_{i,x}^{t+1}\leq\left(1-\frac{\bar{w}}{2}\right)\Delta_{i,x}^{t}+\frac{C_{3}}{(t+1)^{1+v_{x}-v_{z}}},\vspace{-0.2em}$ (17)
with $C_{3}=\frac{2C_{y}\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{0}}{\lambda_{y}^{0}}+\frac{2(C_{1}+d_{z})\sqrt{r}L_{l}\lambda_{x}^{0}}{\lambda_{z}^{0}}+2\sqrt{n_{i}}L_{h}\lambda_{x}^{0}$.
@@
$\vspace{-0.2em}\textstyle\Delta_{i,x}^{t}\!\leq\!\!\frac{C_{x}}{(t+1)^{1+v_{x}-v_{z}}}~{}\text{with}~{}C_{x}\!=\!\frac{4C_{3}}{\bar{w}}\!(\frac{4(1+v_{x}-v_{z})}{e\ln(\frac{4}{4-\bar{w}})}\!)^{1+v_{x}-v_{z}}\!.\vspace{-0.2em}$ (18)
@@
$\vspace{-0.2em}\textstyle\Delta_{i,z}^{t}\leq\frac{C_{z}}{(t+1)^{1+v_{z}}}~{}\text{with}~{}C_{z}=\frac{2C_{4}}{\bar{w}}(\frac{4(1+v_{z})}{e\ln(\frac{2}{2-\bar{w}})})^{1+v_{z}},\vspace{-0.2em}$ (19)
with $C_{4}=C_{x}\sqrt{r}\bar{L}_{h}\lambda_{z}^{0}+2C_{y}\sqrt{r}\bar{L}_{h}\lambda_{z}^{0}(\lambda_{y}^{0})^{-1}+2\sqrt{r}L_{h}\lambda_{z}^{0}$.
@@
$\textstyle\sum_{t=1}^{T}\left(\frac{\Delta_{i,x}^{t}}{\nu_{i,x}}+\frac{\Delta_{i,y}^{t}}{\nu_{i,y}}+\frac{\Delta_{i,z}^{t}}{\nu_{i,z}}\right)\leq\sum_{t=1}^{T}\left(\frac{\sqrt{2}C_{x}}{\sigma_{i,x}(t+1)^{1+v_{x}-v_{z}-\varsigma_{i,x}}}\right.$
$\left.\textstyle\quad+\frac{\sqrt{2}C_{y}}{\sigma_{i,y}(t+1)^{1+v_{y}-\varsigma_{i,y}}}+\frac{\sqrt{2}C_{z}}{\sigma_{i,z}(t+1)^{1+v_{z}-\varsigma_{i,z}}}\right),$ (20)
where $\nu_{i,x}$, $\nu_{i,y}$, and $\nu_{i,z}$ represent Laplace noise variances, which are given in Assumption 4.
(ii) Inequality (20) implies that the cumulative privacy budget $\epsilon_{i}$ is finite even when $T$ tends to infinity since $v_{x}-v_{z}\!>\!\max_{i\in[m]}\{\varsigma_{i,x}\}$, $v_{y}\!>\!\max_{i\in[m]}\{\varsigma_{i,y}\}$, and $v_{z}\!>\!\max_{i\in[m]}\{\varsigma_{i,z}\}$ always hold based on Assumption 4. ∎
We would like to point out that although ensuring accurate convergence, the achievement of rigorous $\epsilon$-LDP of Algorithm 1 does pay a price in convergence rates. We quantify the tradeoff in the following corollary:
@@
For any given cumulative privacy budget $\epsilon_{i}>0,~{}i\in[m]$, the convergence rates of Algorithm 1 are $\mathcal{O}\left(\frac{T^{-\beta}}{\min_{i\in[m]}\{\epsilon_{i}^{2}\}}\right)$ for strongly convex $F(\boldsymbol{x})$ and $\mathcal{O}\left(\frac{T^{-(1-v_{x})}}{\min_{i\in[m]}\{\epsilon_{i}^{2}\}}\right)$ for general convex or nonconvex $F(\boldsymbol{x})$.
@@
(a) Training loss
(b) Test accuracy
(c) Cumulative privacy budget of $1,817$ iterations
Figure 1: Comparison of Algorithm 1 with existing distributed aggregative optimization algorithms in [4] and [21] (under the same cumulative privacy budget), and existing DP approaches in [36] and [48] for distributed optimization, using the “MNIST” dataset. The error bars represent the standard deviation.

### v3

@@ V Privacy analysis
$\textstyle\frac{\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{p})\tilde{z}_{i}^{t}-\nabla l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{\prime p})\tilde{z}_{i}^{\prime t}\|_{1}$ (11)
$\textstyle\leq\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\sum_{p=0}^{t}\|\nabla l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{p})\|_{2}\|\tilde{z}_{i}^{t}-\tilde{z}_{i}^{\prime t}\|_{1}$
$\textstyle\quad+\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\sum_{p\neq k}^{t}\|\nabla l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{p})-\nabla l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{p})\|_{2}\|\tilde{z}_{i}^{\prime t}\|_{2}$
$\textstyle\quad+\frac{\sqrt{n_{i}}\lambda_{x}^{t}}{t+1}\|\nabla l(R_{i}\mathbf{x}_{i}^{t};\xi_{i}^{k})-\nabla l(R_{i}\mathbf{x}_{i}^{\prime t};\xi_{i}^{\prime k})\|_{2}\|\tilde{z}_{i}^{\prime t}\|_{2}$
$\textstyle\leq\frac{\sqrt{n_{i}}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}}(\Delta_{i,z}^{t+1}+\Delta_{i,z}^{t})\!+\!\frac{\sqrt{n_{i}}\bar{L}_{l}d_{i,z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\Delta_{i,x}^{t}\!+\!\frac{2\sqrt{n_{i}}L_{l}d_{i,z}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)},$
where we have used the fact that the convergence of Algorithm 1 ensures $\max_{t>0}\{\|z_{i}^{t}\|_{2}\}\leq d_{i,z}$ for some $d_{i,z}>0$.
Substituting (10) and ($1$) into (9), we arrive at
$\textstyle\Delta_{i,x}^{t+1}\leq\left(1-|w_{ii}|+\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{t}+\frac{\sqrt{n_{i}}\bar{L}_{l}d_{i,z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\right)\Delta_{i,x}^{t}$
@@
$\textstyle\quad+\frac{2\sqrt{n_{i}}L_{h}\lambda_{x}^{t}}{t+1}+\frac{2\sqrt{n_{i}}d_{i,z}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)}.$ (12)
@@
$\|z_{i}^{t+1}-z_{i}^{\prime t+1}\|_{1}\leq(1-|w_{ii}|)\|z_{i}^{t}-z_{i}^{\prime t}\|_{1}$
$\textstyle+\frac{\sqrt{r}\lambda_{z}^{t}}{t+1}\sum_{p\neq k}^{t}\|\nabla_{y}h(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{p})\!-\!\nabla_{y}h(R_{i}\mathbf{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{p})\|_{2}$
$\textstyle+\frac{\sqrt{r}\lambda_{z}^{t}}{t+1}\|\nabla_{y}h(R_{i}\mathbf{x}_{i}^{t},\tilde{y}_{i}^{t};\varphi_{i}^{k})\!-\!\nabla_{y}h(R_{i}\mathbf{x}_{i}^{\prime t},\tilde{y}_{i}^{\prime t};\varphi_{i}^{\prime k})\|_{2}.$
@@
$\Delta_{i,z}^{t+1}\textstyle\leq(1-|w_{ii}|)\Delta_{i,z}^{t}+\sqrt{r}\bar{L}_{h}\lambda_{z}^{t}\Delta_{i,x}^{t}$ (13)
@@
$\textstyle\leq\left(1\!-\!\bar{w}\!+\!\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{t}+\frac{\sqrt{n_{i}}\bar{L}_{l}d_{i,z}\lambda_{x}^{t}}{\lambda_{z}^{t}}\!+\!\sqrt{r}(L_{l}\lambda_{y}^{t}\!+\!\bar{L}_{h}\lambda_{z}^{t})\right)\Delta_{i,x}^{t}$
@@
$\textstyle+\frac{2d_{i,l}\lambda_{y}^{t}}{t+1}+\frac{2\sqrt{n_{i}}L_{h}\lambda_{x}^{t}}{t+1}+\frac{2\sqrt{n_{i}}d_{i,z}L_{l}\lambda_{x}^{t}}{\lambda_{z}^{t}(t+1)}+\frac{2\sqrt{r}L_{h}\lambda_{z}^{t}}{t+1}.$ (14)
Since the decaying rates of stepsizes $\lambda_{x}^{t}$, $\lambda_{y}^{t}$, and $\lambda_{z}^{t}$ satisfy $1\!>\!v_{x}\!>\!v_{z}\!>\!v_{y}\!>\!0$, we can choose proper initial stepsizes such that the following inequality always holds:
$\textstyle\Delta_{i,x}^{t+1}+\Delta_{i,y}^{t+1}+\Delta_{i,z}^{t+1}$ $\textstyle\leq\big(1-\frac{|w_{ii}|}{2}\big)\left(\Delta_{i,x}^{t}+\Delta_{i,y}^{t}+\Delta_{i,z}^{t}\right)$
$\textstyle\quad+\frac{C_{i,0}}{(t+1)^{\min\{1+v_{x}-v_{z},1+v_{y}\}}},$ (15)
with $C_{i,0}\!=\!2(d_{i,l}\lambda_{y}^{0}+\sqrt{n_{i}}\lambda_{x}^{0}(L_{h}+d_{i,z}L_{l}(\lambda_{z}^{0})^{-1})+\sqrt{r}L_{h}\lambda_{z}^{0})$.
By applying Lemma 11 in [15] to (15) (here, in (15), $\Delta_{i,x}^{t}+\Delta_{i,y}^{t}+\Delta_{i,z}^{t}$ corresponds to $\psi_{t}$ in Lemma 11 in [15]) and using the relation $\Delta_{i,x}^{0}\!=\!\Delta_{i,y}^{0}\!=\!\Delta_{i,z}^{0}\!=\!0$, we have
$\Delta_{i,x}^{t}+\Delta_{i,y}^{t}+\Delta_{i,z}^{t}\leq\frac{C_{i,1}}{(t+1)^{\min\{1+v_{x}-v_{z},1+v_{y}\}}},\vskip-1.99997pt$ (16)
with $C_{i,1}\!=\!\frac{4C_{i,0}}{|w_{ii}|}(\frac{4\min\{1+v_{x}-v_{z},1+v_{y}\}}{e\ln(\frac{4}{4-|w_{ii}|})})^{\min\{1+v_{x}-v_{z},1+v_{y}\}}$.
Substituting $\Delta_{i,x}^{t}\leq C_{i,1}(t+1)^{-\min\{1+v_{x}-v_{z},1+v_{y}\}}$ into (8) and using Lemma 11 in [15], we obtain
$\Delta_{i,y}^{t}\leq\frac{C_{i,y}}{(t+1)^{1+v_{y}}}~\text{with}~{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\definecolor{pgfstrokecolor}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}C_{i,y}=\frac{8C_{i,2}}{|w_{ii}|}(\frac{1+v_{y}}{e\ln(\frac{2}{2-|w_{ii}|})})^{1+v_{y}}},\vskip-1.99997pt$ (17)
where $C_{i,2}=(C_{i,1}L_{l}\sqrt{r}+2d_{i,l})\lambda_{y}^{0}$ and $C_{i,1}$ is given in (16).
Substituting $\Delta_{i,z}^{t}\!\leq\!C_{i,1}(t+1)^{-\min\{1+v_{x}-v_{z},1+v_{y}\}}$ and (17) into (12), we have
$\Delta_{i,x}^{t+1}\leq\big(1-\frac{|w_{ii}|}{2}\big)\Delta_{i,x}^{t}+\frac{C_{i,3}}{(t+1)^{1+v_{x}-v_{z}}},\vskip-1.99997pt$ (18)
with $C_{i,3}=\frac{2C_{i,y}\sqrt{n_{i}}\bar{L}_{h}\lambda_{x}^{0}}{\lambda_{y}^{0}}+\frac{2(C_{i,1}+d_{i,z})\sqrt{n_{i}}L_{l}\lambda_{x}^{0}}{\lambda_{z}^{0}}+2\sqrt{n_{i}}L_{h}\lambda_{x}^{0}$.
@@
$\vskip-1.99997pt\textstyle\Delta_{i,x}^{t}\!\leq\!\!\frac{C_{i,x}}{(t+1)^{1+v_{x}-v_{z}}}~\text{with}~{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\definecolor{pgfstrokecolor}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}C_{i,x}\!=\!\frac{16C_{i,3}}{|w_{ii}|}\!(\frac{1+v_{x}-v_{z}}{e\ln(\frac{4}{4-|w_{ii}|})}\!)^{1+v_{x}-v_{z}}}\!.\vskip-1.99997pt$ (19)
@@
$\vskip-1.99997pt\textstyle\Delta_{i,z}^{t}\leq\frac{C_{i,z}}{(t+1)^{1+v_{z}}}~\text{with}~{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\definecolor{pgfstrokecolor}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}\definecolor{XC@mixtmp}{rgb}{0,0,1}C_{i,z}=\frac{8C_{i,4}}{|w_{ii}|}(\frac{1+v_{z}}{e\ln(\frac{2}{2-|w_{ii}|})})^{1+v_{z}}},$ (20)
with $C_{i,4}=C_{i,x}\sqrt{r}\bar{L}_{h}\lambda_{z}^{0}+2C_{i,y}\sqrt{r}\bar{L}_{h}\lambda_{z}^{0}(\lambda_{y}^{0})^{-1}+2\sqrt{r}L_{h}\lambda_{z}^{0}$.
@@
$\textstyle\sum_{t=1}^{T}\left(\frac{\Delta_{i,x}^{t}}{\nu_{i,x}}+\frac{\Delta_{i,y}^{t}}{\nu_{i,y}}+\frac{\Delta_{i,z}^{t}}{\nu_{i,z}}\right)\leq\sum_{t=1}^{T}\left(\frac{\sqrt{2}C_{i,x}}{\sigma_{i,x}(t+1)^{1+v_{x}-v_{z}-\varsigma_{i,x}}}\right$.
$\left.\textstyle\quad+\frac{\sqrt{2}C_{i,y}}{\sigma_{i,y}(t+1)^{1+v_{y}-\varsigma_{i,y}}}+\frac{\sqrt{2}C_{i,z}}{\sigma_{i,z}(t+1)^{1+v_{z}-\varsigma_{i,z}}}\right),$ (21)
where $\nu_{i,x}$, $\nu_{i,y}$, and $\nu_{i,z}$ represent Laplace-noise variances, which are given in Assumption 4.
(ii) Inequality (21) implies that the cumulative privacy budget $\epsilon_{i}$ is finite even when $T\rightarrow\infty$ since $v_{x}-v_{z}>\varsigma_{i,x}$, $v_{y}>\varsigma_{i,y}$, and $v_{z}>\varsigma_{i,z}$ always hold based on Assumption 4.
Note that for any given $\epsilon_{i}>0$, the required initial noise of agent $i$ depends on its weighted degree $|w_{ii}|\!=\sum_{j\in\mathcal{N}_{i}}w_{ij}$. Specifically, according to our analysis in (155), for any given $\epsilon_{i}\!>\!0$, the initial DP-noises of agent $i$ can be chosen as $\sigma_{i,x}\!=\!\frac{3\sqrt{2}C_{x}}{(v_{x}-v_{z}-\varsigma_{i,x})\epsilon_{i}}$, $\sigma_{i,y}\!=\!\frac{3\sqrt{2}C_{y}}{(v_{y}-\varsigma_{i,y})\epsilon_{i}}$, and $\sigma_{i,z}\!=\!\frac{3\sqrt{2}C_{z}}{(v_{z}-\varsigma_{i,z})\epsilon_{i}}$, where $v_{x}$, $v_{y}$, $v_{z}$, $\varsigma_{i,x}$, $\varsigma_{i,y}$, and $\varsigma_{i,z}$ are preset parameters based on Corollary 1, and $C_{i,x}$, $C_{i,y}$, and $C_{i,z}$ are given in (19), (17), and (20), respectively. According to (19), (17), and (20), the constants $C_{i,x}$, $C_{i,y}$, and $C_{i,z}$ are inversely proportional to the weighted degree $|w_{ii}|$. Consequently, a smaller weighted degree implies that agent $i$ requires larger initial noise to ensure a desired cumulative privacy budget $\epsilon_{i}$.
We would like to point out that although ensuring accurate convergence, the achievement of rigorous $\epsilon_{i}$-LDP of Algorithm 1 does pay a price in convergence rates. We quantify the tradeoff in the following corollary:
@@
For any given cumulative privacy budget $\epsilon_{i}>0$ and $i\in[m]$, the convergence rates of Algorithm 1 are $\mathcal{O}\left(\frac{T^{-\beta}}{\min_{i\in[m]}\{\epsilon_{i}^{2}\}}\right)$ for strongly convex $F(\boldsymbol{x})$ and $\mathcal{O}\left(\frac{T^{-(1-v_{x})}}{\min_{i\in[m]}\{\epsilon_{i}^{2}\}}\right)$ for general convex or nonconvex $F(\boldsymbol{x})$.
@@
(a) Comparison of different algorithms
(b) Cumulative privacy budgets
(c) Comparison under different $\delta$
(d) Comparison under different $\lambda$
Figure 1: Experimental results on CNN training using the “MNIST” dataset. (a) and (b) Training loss and cumulative privacy budget comparison of Algorithm 1 with existing distributed aggregative optimization algorithms in [6] and [36] under the same cumulative privacy budget, and existing DP approaches in [26] and [51] for distributed optimization. (c) Training losses of Algorithm 1 under different $\delta$ ($\delta$ is given in Corollary 1). (d) Training losses of Algorithm 1 under different $\lambda$ ($\lambda$ is defined in (22)). The error bars represent the standard deviation.